OpenStudy (anonymous):

Could someone please help me?? I'm stuck. A circle is inscribed in a square. Write and simplify an expression for the ratio of the area of the square to the area of the circle. For a circle inscribed in a square, the diameter of the circle is equal to the side length of the square. (3 points) Show ALL work.

OpenStudy (anonymous):

@ParthKohli @TuringTest Help?

OpenStudy (anonymous):

@ash2326

OpenStudy (ash2326):

Let's first draw a figure Assume that the circle is touching all the sides

OpenStudy (anonymous):

Okay, what comes next?

OpenStudy (ash2326):

Let the side of square be a That's also the diameter of circle What's the are of a square?

OpenStudy (anonymous):

A=s^2 ?

OpenStudy (ash2326):

What's the side of square here? use that in the formula

OpenStudy (anonymous):

A=a^2?

OpenStudy (ash2326):

Good, now what's the radius of circle?

OpenStudy (anonymous):

Ummm, is it A=pi*a^2?

OpenStudy (ash2326):

a is the diameter, so a/2 will be the radius. Now use the formula

OpenStudy (anonymous):

I'm sorry, I'm not sure how to do that.... I'm very bad with this sort of questions.

OpenStudy (ash2326):

No problem, I'll explain you. For a circle inscribed in a square, the diameter of the circle is equal to the side length of the square. Square side is a Diameter of circle will be a Radius=(diameter)/2 \[r=\frac a 2\] Area of circle=\(\pi\times r^2\) \[Area\ of\ circle=\pi\times \frac {a^2}{4}\] Do you get this ?

OpenStudy (anonymous):

Kinda sorta.... Still very confused/:

OpenStudy (ash2326):

Where do you have doubt?

OpenStudy (anonymous):

More like, I don't know where to go from here.

OpenStudy (ash2326):

Ratio of area of square to circle It's easy divide area of square by circle

OpenStudy (anonymous):

Can you take me through it step by step because I'm still not sure how to do that.....

OpenStudy (ash2326):

\[\large \text{Ratio}=\frac{\text{Area of square}}{\text{Area of circle}}=\frac{a^2}{\pi \times \frac {a^2}{4}}\]

OpenStudy (anonymous):

Ok, and where do I go from here?

OpenStudy (ash2326):

Cancel the common terms \[\large \text{Ratio}=\frac{\text{Area of square}}{\text{Area of circle}}=\frac{\cancel {a^2}}{\pi \times \frac {\cancel{a^2}}{4}}=\frac 4 \pi\]

OpenStudy (anonymous):

Ok, what comes next?

OpenStudy (ash2326):

It's over

OpenStudy (anonymous):

Oh! Oh brother.. You see how bad I am at these. Thank you so much